Circular No. 7809

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA OPHIUCHI 2002 A. Retter and S. O’Toole, University of Sydney; R. Stathakis and J.

Pogson, Anglo-Australian Observatory; and T. Naylor, Exeter University, report: “Using the Anglo-Australian Telescope (+ RGO) on Jan. 26.77 UT, we observed the possible nova reported on IAUC 7808. Preliminary analysis of the medium-resolution spectra (400–700 nm) shows that they are dominated by strong emission lines and possibly weak P-Cyg profiles. The strongest lines are Hα, Hβ, and Fe II (multiplets 42, 74, 48, 49, 55, etc.). The FWZI of the Hα and Hβ lines is 2350 ± 100 km/s. The object is, therefore, very likely a classical nova that belongs to the Fe II class, caught in the early decline phase.”

Prediscovery magnitude estimates: Jan. 19.854 UT, 8.9 (T. Seki, Kochi, Japan; photographic; communicated by H. Sato, Sukagawa, Japan, and T. Kato, Kyoto University); 21.854, 8.5 (H. Nishimura, Kakegawa, Shizuoka, Japan; T-Max 400 film; communicated by Y. Muraoka, Hamamatsu, Japan, and Kato); 24.843, 9.0 (Nishimura). Visual magnitude estimates: Jan. 27.333, 9.2 (J. Garcia, Rama Caida, Argentina); 27.658, 8.9 (M. Linnolt, Honolulu, HI); 27.833, 9.0 (A. Pearce, Nedlands, W. Australia).

SUPERNOVA 2002ao IN UGC 9299 P. Martin and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on an unfiltered CCD image taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Jan. 25.5 UT (mag ∼ 14.3). SN 2002ao is located at α = 14^{h}29^{m}35^{s}.74, δ = −0^{o}00′55′′.8 (equinox 2000.0), which is ∼ 16′′ east and 10′′ north of the diffuse nucleus of UGC 9299. The new object was confirmed at mag 14.7 in an image taken with the Beijing Astronomical Observatory (BAO) 0.6-m telescope on Jan. 26.9 by Y. L. Qiu. Images of the same field taken with KAIT on 2001 June 28.3 (limiting mag ∼ 19.0) and with the BAO 0.6-m on 2002 Jan. 4.9 (limiting mag ∼ 17.8) showed nothing at this position. D. West, Mulvane, KS, also obtained confirming CCD images of SN 2002ao with a 0.2-m reflector, yielding position end figures 35^{s}.72, 56′′.5, and the following magnitudes: Jan. 27.492, R_{c} = 14.83; 27.507, V = 14.96.

COMET C/2000 WM1 (LINEAR) Visual m_{1} estimates by A. Pearce, Nedlands, W. Australia (8×40

binoculars): Jan. 1.53 UT, 6.0; 12.53, 6.0; 27.85, 4.6 (outburst).

2002 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green